(b)a body at rest is giving an initial uniform acceleration of 8m/s² for 30s, after which the acceleration is released to 5m/s² for the next 20s.the body maintains this speed attained for 60 seconds,after which it is Brought to rest in 20 seconds 

Draw the velocity time graph of the motion using the information given



(c) using graph, calculate the :

(¡) maximum speed attained during the motion

(¡¡)average retardation of the body is being brought to rest 

(¡¡¡)total distance during the first 50s

(¡v) average speed during the same interval as in (¡¡¡)

(v) total distance travelled by the body 


1
Expert's answer
2021-07-28T17:49:04-0400

The velocity-time graph of the motion 



i) The maximum speed attained during the motion:


"V=160\frac{m}{s}"

ii) The average retardation as the body is being brought to rest:


"a=8\frac{m}{s^2}"

iii) The total distance traveled during the first 50 s:


"s=0.5(160)(20)+(160)(50-20)=6400\ m"

iv) The average speed during the same interval as in (ii):


"v_a=\frac{s}{t}=\frac{6400}{50}=128\frac{m}{s}"

v) The total distance travelled by the body 


"s=0.5(160)(20)+(160)(80-20)\\+0.5(160)(20)=12800\ m"